OpenStudy (james1107):

Which of the following equations has only one solution? x^2 - 8x + 16 = 0 x(x - 1) = 8 x^2 = 16

OpenStudy (dinamix):

the frist one cuz delta =0

OpenStudy (dinamix):

@james1107

OpenStudy (kendricklamar2014):

What is your answer? Can you show me the steps you did to solve each problem?

OpenStudy (dinamix):

and its only

OpenStudy (kendricklamar2014):

@dinamix Direct answers are not allowed on OpenStudy.

OpenStudy (james1107):

thanks cuzzzz

OpenStudy (kendricklamar2014):

\[x^2−8x+16=0\] Step 1: Factor left side of equation. \[(x−4)(x−4)=0\] Step 2: Set factors equal to 0. \[x−4=0 ~ or ~ x−4=0\] \[x=4\]